OpenStudy (anonymous):

My problem is X^2-6x+10=0 and i know the answer is 3+i , 3-i Need step by step process, please thanks.

OpenStudy (ash2326):

@gabz12 do you know quadratic formula?

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (ash2326):

good, we know that quadratic equation is of the following form \[ax^2+bx+c=0\] Our equation is \[x^2-6x+10=0\] Can you plugin the values of a, b and c in the formula you posted.

OpenStudy (anonymous):

\[x=\frac{ -(-6)\sqrt{-6^{2}-4(1)(10)} }{ 2(1)} \] ?

OpenStudy (ash2326):

yes, just one thing. Inside the root it should be (-6)^2 not -(6)^2.

OpenStudy (anonymous):

well that was my mistake then i kept getting \[\frac{ 6\pm \sqrt{-76} }{ 2 }\]

OpenStudy (ash2326):

So now you know the issue

OpenStudy (anonymous):

but now i got \[\frac{ 6\pm \sqrt{4i} }{ 2 }\]

OpenStudy (ash2326):

you can bring out the 4 from the root and cancel the common term from numerator and denominator.

OpenStudy (anonymous):

wouldnt i be left with \[6\pm i\]

OpenStudy (ash2326):

\[\frac{6\pm \sqrt 4 i}{2}\] \[\frac{6\pm 2i}{2}\] divide both numerator terms by 2 and you will get your answer.

OpenStudy (anonymous):

oh i see where my problem was thank you for your time this was very helpful!

OpenStudy (ash2326):

You're welcome :)